# Detecting Spatial Dependence in Transcriptomics Data using Vectorised Persistence Diagrams

### v1

@@ 5.3 Analysing Transcriptomics Data
Having observed that persistence-based approaches give suitable spatial variability tests on simulated data, we next investigate their utility on real datasets as detailed in Section 4.4.
@@ 5.3.1 PH-based methods identify spatially variable genes
First, we decide to focus on one of our proposed methods, namely Betti curves, which performed well across our simulated case study, and investigate spatially variable genes detected by this approach. Then, Fig. 9 shows representative examples of genes that are both the most highly variable and the most spatially variable features reaching the lowest possible p-values in terms of the spatial dependence test. We see that Betti curves identify both genes with low and high expression counts to be spatially variable. For the mouse olfactory bulb (MOB), we see that there is a clear pattern of genes being more highly expressed and clustered in the center of this brain region, such as Mbp, Pcp4, Nrgn, or Gpsm1. What these genes have in common then is that they relate to neural activity in this region of the mouse’s brain. For the beast cancer sample we see some more spares patterns that could indicate disease specific differences in tissue regions and immune response. Finally, for the heart sample the genes COL1A1 and COL1A2 are found to be spatially variable, which are coding for collagen type 1 mainly found in connective tissue and related to wound healing. Overall, studying the set of spatially variable genes in more depth can thus reveal important insights into tissue biology and disease response by specifically focusing on transcripts that vary across tissue regions.
@@ 5.3.2 PH-based methods identify fewer SVGs than Moran’s I
Across the real datasets considered for this study, we report the number of spatially variable genes detected by each method. Results in Fig. 7 then confirm that Moran’s I reports the largest number of significantly spatially auto-correlated genes calling more than $50\%$ of transcripts spatially variable. This is followed by Betti curves, which find more spatially variable patterns for the heart sample, but fewer for the mouse brain and breast cancer sample. Interestingly, across these real datasets, total persistence detects a low number of SVGs, which is comparable to Persistence landscapes, but notably lower than Betti curves. This trend has not been observed across our simulations indicating that the results of using Betti curves or total persistence are not necessarily similar. Overall, the reported numbers of SVGs support our findings from the simulation study in Section 5.1 on the sensitivity and specificity of each approach that lead us to conclude that persistence-based methods identify a smaller number of genes than Moran’s I but do so with a lower type-1 error rate.
@@ 5.3.3 Persistence is robust to the sum of feature values
Next, we investigate the dependence of each score on the total counts measured for each gene as reported in Fig. 7. Ideally, spatial autocorrelation test evaluation also be somewhat robust of whether features have high or low values overall. That is, genes that are found to be spatially variable should not be baised towards genes with high expression levels, but also detect spatial patterns in lowly expressed genes. Using the Spearman correlation with total counts as a sanity check, we see that amongst all methods, Moran’s I generally shows the highest correlation with total counts. This is followed by persistence-based methods which are more robust to the sum of feature values. Finally, sepal is designed to quantify spatial variability independently of total counts, but we find that in some examples it actually correlates more with total counts than persistence-based methods.
@@ 5.3.4 Agreement between methods
Further, Fig. 8 reports the rank correlations between the gene rankings reported by each spatially variable gene detection method. Overall, we once again see a high agreement between Betti curves and total persistence $\rho\in[0.7,0.79]$. This is followed by moderate correlation to persistence landscapes, Moran’s I and SpatialDE with $\rho\in[0.32,0.69]$. Persistence landscapes rank genes in a notably different order than other persistence-based approaches $\rho\in[0.42,0.59]$. Overall, the least agreement is reported between persistence-based methods and sepal with $\rho\in[0.24,0.39]$. Overall, our results point towards a relatively high disagreement between methods. Similar to the simulation study, we find that persistence summarises distinct information on spatial patterns not encoded by existing SVG detection methods.
@@ 6 Discussion
Throughout our study, we proposed randomised permutation testing procedures for spatial variability in omics data via Betti curves or persistence landscapes. We investigated these tests in comparison to baseline methods Moran’I, sepal, SpatialDE and total persistence Byers et al. [11].
@@
On simulated data, we showed that persistence is more robust to zero-inflation than Moran’s I showing higher AUPCR and a better detection rate of truly variable features amongst the highest ranked genes. At a fixed significance cut-off we further find that persistence-based methods lead to slightly lower sensitivity but higher specificity making us conclude that persistence is potentially more useful than Moran’s I in applications that require more conservative estimates. We thus demonstrate in-silico that persistence curves offer stable approaches to quantify spatial dependence in feature values.
On real data, we further observe that persistence is less dependent on the sum of feature values than Moran’s I indicating that it summarises spatial patterns in a more robust manner. Finally, we visualise genes with representative spatial patterns that are both spatially and highly variable and describe their qualitative biological relevance. Thus, we describe the practical benefits of using persistence-based approaches over Moran’s I and showcase examples where persistence curves are superior to baseline methods.
@@
We find that a non-parametric approach is less reliant on specific distributional assumptions about the spatial distribution of feature values, in contrast to parametric methods such as SpatialDE, making our approach preferable when such assumptions are not satisfied [30], which arguably remains a contested question for omics data in particular. While our analysis focuses on illustrating examples on the use of persistent homology in the domain of spatial omics data analysis, our methodology can be generalised to other types of spatial data, such as spatial graphs with labelled vertices or edges. We thus propose a flexible method for detecting spatially variable features based on application-driven filtrations.
@@
However, we want to also note the limitations of our study. The computational costs of non-parametric methods, especially the computation-heavy nature of permutation testing, remains a concern in practice [30]. Computing a persistence curve for one spatial feature itself is extremely fast taking less than one second on a local CPU. However, the proposed spatial variablity tests pose some scalability constraints. Permutation testing requires thousands of repetitions resulting in a notable increase in computational costs [24]. While per feature this test might then take less than a minute, this scaling behaviour does not suffice when the goal is to study hundreds of thousands of genes simultaneously. Sequential permutation testing could be employed to reduce some of these costs [19], but might not lead to sufficient computational improvements overall. There thus exists a great motivation for further work to develop alternative statistical tests using potentially parametric results, such as ones that could be derived from a universal null distribution for persistence diagrams [8] to speed up the proposed testing procedures.
@@
In terms of the way persistence is used throughout this study, we investigated zeroth dimensional persistent homology because of its computational efficiency as well as the utility of encoding local peaks for detecting spatial patterns. However, we note that one could follow an approach similar to Hickok et al. [21], extend the filtration by its exterior and compute first dimensional persistent homology. This would allow to deduce information about the geographical location of local minima via tracking the generators of 1D homology classes, which could be combined with statistical methods to quantify exactly how spatial patterns vary from spatial randomness.

### v2

@@ 5.3 Analysing Transcriptomics Data
Having observed that persistence-based approaches give suitable spatial variability tests on simulated data, we next investigate their utility on real datasets of spatial tissue samples as detailed in Section 4.4.
@@ 5.3.1 PH-based methods identify spatially variable genes
First, we decide to focus on one of our proposed methods, namely Betti curves, which performed well across our simulated case study, and investigate spatially variable genes detected by this approach. Fig. 9 shows representative examples of genes that are both the most highly variable and the most spatially variable features reaching the lowest possible p-values in terms of the spatial dependence test. We see that Betti curves identify both genes with low and high expression counts to be spatially variable. For the mouse olfactory bulb (MOB), we see that there is a clear pattern of genes being more highly expressed and clustered in the centre of this brain region, such as Mbp, Pcp4, Nrgn, or Gpsm1. What these genes have in common is that they relate to neural activity in this region of the mouse’s brain [24]. For the breast cancer sample, we see some more sparse patterns that could indicate disease-specific differences in tissue regions and immune response. Finally, for the heart sample, the genes COL1A1 and COL1A2 are found to be spatially variable, which are coding for collagen type 1, mainly found in connective tissue and related to wound healing. These genes have been found to be unregulated in tissue regions impacted by myocardial infarction [44] and are thus of biological relevance for understanding the spatial composition of this tissue sample. Overall, studying the set of spatially variable genes in more depth can thus reveal important insights into tissue biology and disease response by specifically focusing on transcripts that vary across tissue regions.
@@ 5.3.2 PH-based methods identify fewer SVGs than Moran’s I
Across the real datasets considered for this study, we report the number of spatially variable genes detected by each method. Results in Fig. 7 then confirm that Moran’s I reports the largest number of significantly spatially autocorrelated genes, calling more than $50\%$ of transcripts spatially variable. This is followed by Betti curves, which find more spatially variable patterns for the heart sample, but fewer for the mouse brain and breast cancer sample. Interestingly, across these real datasets, total persistence detects a low number of SVGs, which is comparable to persistence landscapes but notably lower than Betti curves. This trend has not been observed across our simulations, indicating that the results of using Betti curves or total persistence are not necessarily similar. Overall, the reported numbers of SVGs support our findings from the simulation study in Section 5.1 on the sensitivity and specificity of each approach. We conclude that persistence-based methods identify a smaller number of genes than Moran’s I, but do so with a lower type-1 error rate.
@@ 5.3.3 Persistence is robust to the sum of feature values
Next, we investigate the dependence of each score on the total counts measured for each gene as reported in Fig. 7. Ideally, spatial dependence tests should be robust to whether features have high or low values overall. That is, a good method for detecting spatially variable genes should not be biased towards genes with high expression levels, but also detect spatial patterns in lowly expressed genes. We report the Spearman correlation with total counts as a sanity check. Amongst all methods, we see that Moran’s I generally shows the highest correlation with total counts. In contrast, persistence-based methods are more robust to the sum of feature values. Finally, sepal is designed to quantify spatial variability independently of total counts, but we find that in some examples it actually correlates more with total counts than persistence-based methods.
@@ 5.3.4 Agreement between methods
Further, Fig. 8 reports the rank correlations between the gene rankings reported by each spatially variable gene detection method. Overall, we once again see a high agreement between Betti curves and total persistence, with $\rho\in[0.7,0.79]$. This is followed by moderate correlation to persistence landscapes, Moran’s I, and SpatialDE, with $\rho\in[0.32,0.69]$. Persistence landscapes rank genes in a notably different order than other persistence-based approaches, with $\rho\in[0.42,0.59]$. Overall, the least agreement is reported between persistence-based methods and sepal, with $\rho\in[0.24,0.39]$. Overall, our results point towards a relatively high disagreement between methods. Similar to the simulation study reported in Fig. 5, we find that persistence summarises distinct information on spatial patterns not encoded by existing SVG detection methods.
@@ 6 Discussion
Throughout our study, we proposed randomised permutation testing procedures for spatial variability in omics data via Betti curves or persistence landscapes. We investigated these tests in comparison to baseline methods Moran’s I, sepal, SpatialDE, and total persistence.
@@
On simulated data, we showed that persistence is more robust to zero-inflation than Moran’s I, showing higher AUPRC and a better detection rate of truly variable features amongst the highest ranked genes. At a fixed significance cut-off, we further find that persistence-based methods lead to slightly lower sensitivity but higher specificity, suggesting that persistence is potentially more useful than Moran’s I in applications that require more conservative estimates. We thus demonstrate in-silico that persistence curves offer stable approaches to quantify spatial dependence in feature values. On real data, we further observe that persistence is less dependent on the sum of feature values than Moran’s I, indicating that it summarises spatial patterns in a more robust manner. Finally, we visualise genes with representative spatial patterns that are both spatially and highly variable and describe their qualitative biological relevance. Thus, we describe the practical benefits of using persistence-based approaches over Moran’s I and showcase examples where persistence curves are superior to baseline methods.
@@
Overall, we find that a non-parametric approach is less reliant on specific distributional assumptions about the spatial distribution of feature values, in contrast to parametric methods such as SpatialDE. This makes our approach preferable when such assumptions are not satisfied [31], which arguably remains a contested question for omics data in general. While our analysis focuses on illustrating examples on the use of persistent homology in the domain of spatial omics data analysis, our methodology can be generalised to other types of spatial data, such as spatial graphs with labelled vertices or edges. We thus propose a flexible method for detecting spatially variable features based on application-driven filtrations.
@@
However, we want to also note the limitations of our study. The computational costs of non-parametric methods, especially the computation-heavy nature of permutation testing, remains a concern in practice [31]. Computing a persistence curve for one spatial feature itself is extremely fast, taking less than one second on a local CPU. However, the proposed spatial variability tests pose some scalability constraints. Permutation testing requires thousands of repetitions, resulting in a notable increase in computational costs [25]. While per feature this test might then take less than a minute, this scaling behaviour does not suffice when the goal is to study hundreds of thousands of genes simultaneously. Sequential permutation testing could be employed to reduce some of these costs [19], but might not lead to sufficient computational improvements overall. There thus exists a great motivation for further work to develop alternative statistical tests using potentially parametric results, such as ones that could be derived from a universal null distribution for persistence diagrams [8] to speed up the proposed testing procedures.
@@
In terms of the way persistence is used throughout this study, we investigated zero-dimensional persistent homology because of its computational efficiency, as well as the utility of encoding local peaks for detecting spatial patterns. However, we note that one could follow an approach similar to Hickok et al. [21] to extend the filtration by its exterior and compute one-dimensional persistent homology. This would permit deducing information about the geographical location of local minima via tracking the generators of one-dimensional homology classes, which could be combined with statistical methods to quantify exactly how spatial patterns vary from spatial randomness.
Acknowledgments
The authors are grateful for the constructive and detailed feedback given by the reviewers. Their questions and feedback greatly helped in revising this manuscript. We thank the reviewers for believing in the merits of this work. We further want to thank Mairi McClean, Michaela Müller, Emily Simons, and Jera Toporiš for their help in proofreading the paper. K.L. is supported by the Helmholtz Association under the joint research school ‘Munich School for Data Science (MUDS).’ B.R. was partially supported by the Bavarian state government with funds from the Hightech Agenda Bavaria. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI).